OpenStudy (anonymous):

what is the slope of the line with equation y-5= 3/5 (x-10)

OpenStudy (anonymous):

m represents slope which will be 3/5.

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

@Stephanie1000000

OpenStudy (anonymous):

so it would be y-y1= 3/5 (x-x1) but i dont know where to go from there

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

Are you trying to put the equation into standard form?

OpenStudy (anonymous):

well im trying to figure out the slope so i should put it in standard form?

OpenStudy (anonymous):

Then the slope would be 3/5 as shown in your equation.